OpenStudy (anonymous):

rachels brother is 33 years old . he is 3 times rachels age 16 years ago . how old is rachel ?

OpenStudy (anonymous):

its multiple choice so the answers could either be 11 , 27 , 16, or 24

Parth (parthkohli):

\[\rm b = 33 \wedge b = 3(r - 16) \implies 33=3(r - 16)\]

Parth (parthkohli):

That speaks - if b = 33 and b also = 3(r - 16), then 33 = 3(r - 16).

OpenStudy (anonymous):

i dont really need the work to show ho you did it , the answer by it self is fine :)

OpenStudy (jiteshmeghwal9):

Let rachel's age be n 16 years ago Rachel will be n-16. 3(n-16)=33

Parth (parthkohli):

\[\rm {33 \over 3} = r - 16\]\[{33 \over 3} + 16 = r\]

OpenStudy (anonymous):

thank you

Parth (parthkohli):

You're welcome!

OpenStudy (jiteshmeghwal9):

as well as my equation was also right :P